how many outcomes are possible for a car with 2 models, 8 colors, and any combination of 4 options?

Found 4 solutions by ewatrrr, ikleyn, mccravyedwin, greenestamps:
Answer by ewatrrr(24785)

 
Hi,
Model I has 8 colors & 4 options, Model II has 8 colors & 4 options
Model I outcomes possible:
8(4C4 + 4C3 + 4C2 + 4C1 ) = 8(1 + 4 + 6 + 4) = 8(15)
Same number of Outcomes for Model II
Total outcomes: 2(8)(15)
Wish You the Best in your Studies.

Answer by ikleyn(53763)

.

As worded, this post MAKES no SENSE,

since it does not define the notion/conception of "options".


Good for re-cycling, only.



Since the post is NONSENSE, do not take seriously that solution, which follows.

It MAKES no SENSE, too.


/\/\/\/\/\/\/


Post - note after reading the post by @greenestamps.


            Dear @greenestamps, in opposite to your opinion,  I  PERFECTLY  understand and understood the meaning of the problem.

            Nevertheless,  I think that in such form no one post may come to the forum and no one post can be accepted.

            According to it,  my reaction was.


            @greenestamps,  PLEASE  NEVER  think about me that I have a lack of brain cells in my head.

            I have them  quite  ENOUGH.



Answer by mccravyedwin(421)

Be sure to post the COMPLETE problem.  You did not give us the number of
possible available options that we can select 4 from.  Without that number the
problem cannot be worked.

Edwin

Answer by greenestamps(13334)



Two tutors provided responses showing that they did not understand the problem as presented; another tutor provided a response that I think reads something into the problem that is not there.

With 2 models and 8 colors, there are 2*8=16 combinations of model and color.

Then "any combination of 4 options" means there are 4 options available, and you can choose ANY combination of them, including none of them.

Counting the number of combinations of 4 of the options is like finding the number of subsets of a set containing 4 elements; including the empty set (choosing none of the 4 options). The number of subsets of a set with n elements is 2^n, so the number of possible combinations of the 4 options is 2^4=16.

So the total number of possible cars is 2*8*16 = 256.